# On the contraction properties of Sinkhorn semigroups

### v1

@@ Abstract
We develop a novel semigroup contraction analysis based on Lyapunov techniques to prove the exponential convergence of Sinkhorn equations on weighted Banach spaces. This operator-theoretic framework yields exponential decays of Sinkhorn iterates towards Schrödinger bridges with respect to general classes of $\phi$-divergences as well as in weighted Banach spaces. To the best of our knowledge, these are the first results of this type in the literature on entropic transport and the Sinkhorn algorithm. We also illustrate the impact of these results in the context of multivariate linear Gaussian models as well as statistical finite mixture models including Gaussian-kernel density estimation of complex data distributions arising in generative models.
Keywords: Entropic Optimal Transport, Sinkhorn semigroups, Schrödinger bridges, contraction coefficients, Lyapunov functions.
@@ 1.1 Background
The Schrödinger bridge problem is to find a random evolution with minimal relative entropy w.r.t. some reference physical model and with prescribed distributions of the initial and final time random states. This convex minimization problem is a regularized version of the Monge-Kantorovich optimal transport with a relative entropy constraint, see  [56, 66, 68] for recent surveys with numerous references on this subject.
@@
These optimal transport methodologies have become state-of-the-art tools in a variety of application domains, including generative modeling and machine learning [5, 31, 54, 68], statistical barycenter problems [1, 4, 10, 24, 29, 32], economy [13], computer vision [42, 76], control theory [18, 19], and many others.
@@
Under mild integrability conditions, the Schrödinger bridge problem has a unique solution which can be expressed as the fixed point of Sinkhorn equations also referred to as the iterative proportional fitting procedure. For a more thorough discussion on the origins of Sinkhorn iterations, we refer to the early and pioneering articles [39, 46, 55, 74, 75] as well as the more recent studies [18, 28, 57].
@@
Sinkhorn iterations are based on nonlinear integral transformations over the whole state space. Exact finite-dimensional solutions cannot be computed for general models. However, for finite state spaces the Sinkhorn algorithm is a well known simple procedure that repeatedly applies column and row normalization on matrix spaces [39, 46, 55, 74, 75]. For Gaussian models, Sinkhorn equations can also be solved analytically using sequentially Bayes conjugacy principles [2, 24, 25, 51].
@@
Most of the literature on the exponential convergence of Sinkhorn equations is concerned with finite state spaces [14, 45, 74, 75, 77] as well as compact state spaces or bounded cost functions using Hilbert projective metrics techniques [18, 33, 47, 40]. The recent article [2] also presents a refined convergence analysis for Gaussian models based on matrix Riccati difference equations. It is clearly out of the scope of this article to review all the contributions in this field and we simply refer to the recent book [68] as well as the more recent articles [2, 20, 23] and the references therein.
@@
Sub-linear convergence rates of Sinkhorn equations have also been developed in the articles [3, 17, 44]. Linear rates on non-necessarily compact spaces were first obtained by Léger in [58] using elegant gradient descent and Bregman divergence techniques (see also the recent articles [31, 52]). Refined convergence rates at least one order faster for a general class of models including sub-Gaussian target marginals has also been developed in [48] under some exponential moments estimates on Schrödinger potentials.
@@
There are very few articles on the exponential convergence of Sinkhorn iterates for general models on non-compact spaces and unbounded cost functions that apply to Gaussian models, with the notable exception of two recent significant contributions [23, 20]. In [23], the authors investigate quantitative contraction rates for target marginal distributions with an asymptotically positive log-concavity profile and quadratic cost functions associated with a sufficiently large regularization parameter. These exponential entropic decays were recently refined in [20] to apply to all values of the regularization parameter.
To the best of our knowledge, the exponential convergence of Sinkhorn iterates in the context of statistical mixture models remains an important open research question.
@@
As underlined in [2], Sinkhorn semigroups belong to the class of time varying Markov processes sharing a common invariant target measure at each individual step. These classes of time varying Markov semigroups arise in a variety of areas in applied mathematics including on nonlinear filtering, physics and molecular chemistry, see for instance [6, 34, 35, 36, 70, 71, 72] and references therein.
@@
The stability analysis of time inhomogenous Markov processes is a broad and notoriously difficult subject. The articles [43, 70, 71] provide several successful probabilistic and analytical tools, including coupling methods, spectral analysis as well as functional inequalities such as Nash or log-Sobolev inequalities. Sinkhorn semigroups being defined sequentially by nonlinear conjugate transformations, the application of these refined analytical techniques in this context is rather challenging. There is a need for new tools in order to handle these models.
1.2 Contribution
In this paper, we introduce a novel approach to Sinkhorn exponential convergence analysis that relies on Lyapunov techniques and contraction coefficients on weighted Banach spaces.
@@ 1.2 Contribution
To be specific, we investigate Lyapunov approaches and related weighted-norm contraction techniques to analyze the convergence of Sinkhorn semigroups. These semigroup techniques are basic quantitative methods for studying the stability of nonlinear and time-varying stochastic models, see for instance [36] as well as Section 2 in [6]. In the context of time-homogeneous models, we also refer to the pioneering articles by Hasminskii [49] and Meyn and Tweedie [60], see also the book [61] and references therein. These powerful off-the-shelf semigroup techniques are mainly based on two conditions:
  • 1)
    the existence of a Lyapunov function satisfying a drift-type condition, and
  • 2)
    a local minorization condition on the sub-level sets of the Lyapunov function.
@@
Such drift-minorization approaches are well-established techniques to localize the minorization conditions on the compact sub-level sets of the Lyapunov function. In this context, Markov operators act on weighted Banach spaces and operator norms coincide with the contraction coefficients of the Markov integral operator.
Based on these techniques, the analysis developed in this article yields novel contraction estimates with respect to general divergence criteria. It also applies to multivariate Gaussian models as well as statistical finite mixture models.
1.3 Outline of the paper
@@
After presenting some basic notation in Section 1.4, we begin with a brief review on Schrödinger bridge problems and Sinkhorn semigroups in Section 2 (and we refer to [2] for a more detailed discussion on these models within our framework). Some background on Lyapunov approaches and weighted-norm contraction properties of Markov integral operators is provided in Section 3, with additional details in Appendix C. Section 3.1 recalls some basics facts on the Hilbert projective metric, operator norms, entropy and related $\phi$-divergences. The operator contraction coefficients associated with these criteria are introduced in Section 3.2.
@@
With these notations at hand, Section 4 shows how to adapt this operator-theoretic framework to the convergence analysis of Sinkhorn semigroups. Section 4.1 presents some general drift-minorization conditions that enable us to estimate the contraction coefficients of Sinkhorn semigroups. The main contraction theorems introduced in this article are presented in Section 4.2 and Section 4.3.
We illustrate the impact of these results in Section 5, which provides several examples, including strong mixing models, Gaussian models as well as Schödinger bridges between finite mixture distributions.
@@
Section 5.2 discusses the contraction properties of Gaussian Sinkhorn models. Last but not least, the convergence analysis of Sinkhorn semigroups in the context of statistical finite mixture models, including Gaussian mixture, is carried out in Section 5.3.
Most of the technical proofs, as well as some extended analyses, are provided in Appendices A through E.
@@ 1.4 Some basic notation
Let ${\cal M}(\mathbb{X})$ be the set of locally bounded signed measures on $\mathbb{X}$, and let ${\cal M}_{b}(\mathbb{X})\subset{\cal M}(\mathbb{X})$ the subset of bounded measures. Let ${\cal M}_{1}(\mathbb{X})\subset{\cal M}_{b}(\mathbb{X})$ be the convex subset of probability measures on $\mathbb{X}$. The Lebesgue integral of a integrable function $f\in{\cal B}(\mathbb{X})$ with respect to (w.r.t.) some $\nu\in{\cal M}(\mathbb{X})$ is denoted by
$\nu(f):=\int_{\mathbb{X}}f(x)~{}\nu(dx)$
For indicator functions $f=1_{A}$ sometimes we write $\nu(A)$ instead of $\nu(1_{A})$.
@@
$(\mu{\cal K})(dy):=\int~{}\mu(dx){\cal K}(x,dy)\quad\mbox{\rm and}\quad{\cal K}(f)(x):=\int{\cal K}(x,dy)f(y)$ (1)
@@
Given a Markov transition ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and a measure $\mu\in{\cal M}(\mathbb{X})$ we see that
$(\mu\times{\cal K})(d(x,y)):=\mu(dx)\,{\cal K}(x,dy)\quad\mbox{\rm and}\quad(\mu\times{\cal K})^{\flat}(d(x,y))=\mu(dy)\,{\cal K}(y,dx).$
@@
2 Description of the models
2.1 Entropic optimal transport
$\lambda_{U}(dx):=e^{-U(x)}~{}\lambda(dx)\quad\mbox{\rm and}\quad\nu_{V}(dy):=e^{-V(y)}~{}\nu(dy)$
In the latter expression, $W(x,y)$ stands for some potential function on $(\mathbb{X}\times\mathbb{Y})$. In optimal transport literature $W$ is sometimes called the cost function.
Consider the reference probability measure
${\cal P}_{0}(d(x,y)):=(\lambda_{U}\times{\cal Q})(d(x,y)):=\lambda_{U}(dx)~{}{\cal Q}(x,dy).$
The entropic transport problem (a.k.a. dynamic Schrödinger bridge problem) w.r.t. the reference measure ${\cal P}_{0}$ is formulated as
$\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\cal P}~{}|~{}{\cal P}_{0}),$ (3)
where $\text{Ent}({\cal P}~{}|~{}{\cal P}_{0})=\int\log\frac{\delta{\cal P}}{d{\cal P}_{0}}d{\cal P}$ is the relative entropy (or Kullback-Leibler divergence) between ${\cal P}$ and ${\cal P}_{0}$. It is implicitly assumed there exists some ${\cal P}\in\Pi_{U,V}$ such that $\mbox{\rm Ent}({\cal P}~{}|~{}{\cal P}_{0})<\infty$. This condition ensures the existence of a Schrödinger bridge distribution $\mathbb{P}$ solving (3) (cf. the seminal article by Csiszár [27], as well as Section 6 in the Lecture Notes by Nutz [66], see also the survey article by Léonard [56] and references therein).
2.2 Schrödinger bridges
The Schrödinger bridge distribution $\mathbb{P}$ defined in (3) can rarely be computed analytically. However, solutions can be approximated efficiently using the Sinkhorn algorithm, also referred to as the iterative proportional fitting procedure [28, 74, 75]. This procedure is defined sequentially for any $n\geq 0$ by a collection of probability distributions

### v2

@@ Abstract
We develop a novel semigroup stability analysis based on Lyapunov techniques and contraction coefficients to prove exponential convergence of Sinkhorn equations on weighted Banach spaces. This operator-theoretic framework yields exponential decays of Sinkhorn iterates towards Schrödinger bridges with respect to general classes of $\phi$-divergences and Kantorovich-type criteria, including the relative entropy, squared Hellinger integrals, $\alpha$-divergences as well as weighted total variation norms and Wasserstein distances. To the best of our knowledge, these contraction inequalities are the first results of this type in the literature on entropic transport and the Sinkhorn algorithm.
We also provide Lyapunov contractions principles under minimal regularity conditions that allow to provide quantitative exponential stability estimates for a large class of Sinkhorn semigroups. We apply this novel framework in a variety of situations, ranging from polynomial growth potentials and heavy tailed marginals on general normed spaces to more sophisticated boundary state space models, including semi-circle transitions, Beta, Weibull, exponential marginals as well as semi-compact models. Last but not least, our approach also allows to consider statistical finite mixture of the above models, including kernel-type density estimators of complex data distributions arising in generative modeling.
Keywords: Entropic Optimal Transport, Sinkhorn semigroups, Schrödinger bridges, contraction coefficients, weighted total variation norms, Lyapunov functions, Kantorovich and entropy criteria, Wasserstein semi-distances.
@@ 1 Introduction
1.1 Entropic optimal transport
$\lambda_{U}(dx):=e^{-U(x)}~\lambda(dx)\quad\mbox{\rm and}\quad\nu_{V}(dy):=e^{-V(y)}~\nu(dy)$ (1)
In the latter expression, the transition potential $W(x,y)$ stands for some real valued function on $(\mathbb{X}\times\mathbb{Y})$. Consider the reference probability measure constructed as
${\cal P}_{0}(d(x,y)):=(\lambda_{U}\times{\cal Q})(d(x,y)):=\lambda_{U}(dx)~{\cal Q}(x,dy).$ (3)
The entropic transport problem (or Schrödinger bridge problem) with respect to (w.r.t.) the reference measure ${\cal P}_{0}$ is formulated as
$\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\cal P}~|~{\cal P}_{0}),\quad\mbox{\rm with the relative entropy}\quad\text{Ent}({\cal P}~|~{\cal P}_{0}):=\int\log\frac{d{\cal P}}{d{\cal P}_{0}}~d{\cal P}.$ (4)
It is implicitly assumed there exists some ${\cal P}\in\Pi_{U,V}$ such that $\mbox{\rm Ent}({\cal P}~|~{\cal P}_{0})<\infty$. This condition ensures the existence of a Schrödinger bridge distribution $\mathbb{P}$ solving (4) (cf. the seminal article by Csiszár [csiszar-2], as well as Section 6 in the Lecture Notes by Nutz [nutz], and also the survey article by Léonard [leonard] and references therein). Let us note that the product distribution
$(\lambda_{U}\otimes\nu_{V})(d(x,y)):=\lambda_{U}(dx)~\nu_{V}(dy)$
trivially belongs to $\Pi_{U,V}$. In addition, choosing ${\cal P}=(\lambda_{U}\times{\cal K})\,\in\,\Pi_{U,V}$ for some Markov transition ${\cal K}(x,dy)$ we have
$\frac{d{\cal P}}{d{\cal P}_{0}}(x,y)=\frac{d(\lambda_{U}\otimes\nu_{V})}{d{\cal P}_{0}}(x,y)\times\frac{d{\cal P}}{d(\lambda_{U}\otimes\nu_{V})}(x,y)$
which yields
$\mbox{\rm Ent}({\cal P}~|~{\cal P}_{0})=\int~W(x,y)~{\cal P}(d(x,y))+\mbox{\rm Ent}({\cal P}~|~\lambda_{U}\otimes\nu_{V})-\int V(y)\nu_{V}(dy).$
Choosing a potential transition of the form $W(x,y)={\cal W}(x,y)/t$, for some parameter $t>0$, problem (4) is equivalent to the optimal entropic transport problem
$\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\left(\int~{\cal W}(x,y)~{\cal P}(d(x,y))+t~\mbox{\rm Ent}({\cal P}~|~\lambda_{U}\otimes\nu_{V})\right).$ (5)
In the optimal transport literature, function $\cal W$} is sometimes called the cost function and the parameter $t>0$ is interpreted as a regularization parameter. In this context, the non-regularized problem corresponding to $t=0$ coincides with the Kantorovich optimal transport problem with cost function $\cal W$}. For a more detailed discussion on this subject we refer to [dobrushin2006perturbation, rachev1985monge, villani2008optimal] and the numerous references therein.
@@ 1.2 Background
The Schrödinger bridge problem (4) is to find a random evolution with minimal relative entropy w.r.t. some reference physical model and with prescribed distributions of the initial and final time random states. This convex minimization problem is a regularized version of the Monge-Kantorovich optimal transport with a relative entropy constraint, see  [leonard, nutz, peyre] for recent surveys with numerous references on this subject.
@@
Optimal transport methodologies have become state-of-the-art tools in a variety of application domains, including generative modeling and machine learning [arjovsky, doucet-bortoli, kolouri, peyre, rioux2023semi, rioux2024entropic], statistical barycenter problems [agueh, andersen, bigot, cramer-2, cuturi-w, loubes], economy [bojitov], computer vision [dominitz, solomon], control theory [chen, chen-2], and many others.
@@
Under mild integrability conditions, the Schrödinger bridge problem has a unique solution which can be expressed as the fixed point of Sinkhorn equations also referred to as the iterative proportional fitting procedure. For a more thorough discussion on the origins of Sinkhorn iterations, we refer to the early and pioneering articles [deming, fortet, kruithof, sinkhorn-2, sinkhorn-3] as well as the more recent studies [chen, cuturi, leonard-2].
@@
Sinkhorn bridges are based on nonlinear integral transformations over the whole state space (cf. Section 2) and exact finite-dimensional solutions cannot be computed for general models. However, for finite state spaces the Sinkhorn algorithm is a well known simple procedure that repeatedly applies column and row normalization on matrix spaces [deming, fortet, kruithof, sinkhorn-2, sinkhorn-3]. For Gaussian models, Sinkhorn equations can also be solved analytically using sequentially Bayes conjugacy principles [adm-24, cramer-2, cramer-3, janadi].
@@
Most of the literature on the exponential convergence of Sinkhorn equations is concerned with finite state spaces [borwein, fienberg, sinkhorn-2, sinkhorn-3, soules] as well as compact state spaces or bounded cost functions using Hilbert projective metrics techniques [chen, deligiannidis, franklin, marino]. See [greco2023non] for a refined analysis of Sinkhorn iterates and their gradient on the $d$-dimensional torus as well as the more recent article [chizat2025sharper] on compact spaces and bounded cost functions.
As remarked in Section 3 of [cohen], the Hilbert metric dominates all $\phi$-divergences, thus we expect rather strong conditions to ensure the contraction of Sinkhorn transitions w.r.t. the Hilbert metric, see for instance [borwein, chen, deligiannidis, franklin] and Section 4.1 in the present article. Contraction estimates w.r.t. the Hilbert metric yield exponential decays w.r.t. $\phi$-divergences, including the total variation norm, the relative entropy and Wasserstein distances of any order. Nevertheless they can rarely be used to derive contraction estimates with respect to some $\phi$-divergence criteria.
At this point, we underline that even on totally bounded state spaces such as the open unit interval $\mathbb{X}=]0,1[=\mathbb{Y}$ the cost function $W$ as well as the potential functions $(U,V)$ may be unbounded and the reference Markov transition may have a unit Birkhoff contraction coefficient (i.e. equal to $1$), see for instance Section 3.2.1 and the series of illustrations discussed in Section 5.3 and Section 5.4.
The landmark example of quadratic costs on Euclidian spaces is another class of unbounded costs functions. In this context, the recent article [adm-24] also presents a refined stability analysis of Sinkhorn bridges for Gaussian models based on matrix Riccati difference equations. It is clearly out of the scope of this article to review all the contributions in this field and we simply refer to the recent book [peyre] as well as the more recent articles [adm-24, chiarini, durmus] and the references therein.
@@
Sub-linear convergence rates of Sinkhorn equations have also been developed in the articles [alts-2017, chak-2018, dvu-2017]. Linear rates on non-necessarily compact spaces were first obtained by Léger in [leger] using elegant gradient descent and Bregman divergence techniques (see also the recent articles [doucet-bortoli, karimi-2024]). Refined convergence rates at least one order faster for a general class of models including sub-Gaussian target marginals has also been developed in [promit-2022] under some exponential moments estimates on Schrödinger potentials.
@@
There are very few and recent articles on the exponential convergence of Sinkhorn iterates for general models on non-compact spaces and unbounded cost functions that apply to Gaussian models, with the notable exception of three significant contributions [chiarini, durmus, eckstein2025hilbert].
The first exponential stability analysis of unbounded models and non necessarily compactly supported marginals is presented in the article [durmus]. In this study, the authors investigate quantitative contraction rates for target marginal distributions with an asymptotically positive log-concavity profile and quadratic symmetric cost functions associated with a sufficiently large regularization parameter. In our framework this particular class of cost functions corresponds to symmetric Gaussian reference transitions.
The approach taken in [durmus] relies on well known links between gradient and Hessian of Sinkhorn and Schrödinger potentials along Hamilton-Jacobi-Bellman equations with conditional expectations and covariances [chewi2023entropic, pooladian2021entropic], see also [del2025entropic]. The main conditions in [durmus] are expressed in terms of the integrated convexity profiles of marginal potentials. Extending ideas from [conforti2024weak] the proofs are based on the propagation of these convexity profiles based on diffusion reflection techniques introduced in [eberle2016reflection].
The exponential entropic decays presented in [durmus] were further developed and recently refined in [chiarini] to apply to all values of the regularization parameter and some classes of perturbation of log-concave models. The article [chiarini] has opened new perspectives in the convergence analysis for entropically regularized transport problems with quadratic symmetric costs. In the context of distance-like cost functions, one of the basic tool is a remarkable stability theorem, Theorem 1.1 in [chiarini], that allows to control the entropy of Schrödinger bridges by the entropy of their marginal distributions. However this stability approach requires to estimate the semi-concavity properties of generally unknown dual Schrödinger bridge transitions and it only applies to entropically regularized transport problems with symmetric quadratic-type costs. Thus, this framework excludes the important cases when the reference transition is associated with linear Gaussian transitions arising in Ornstein-Uhlenbeck-type diffusion generative models and denoising diffusions.
In this connection, the more recent article [del2025stability] complements the stability approach in  [chiarini] with a novel entropic continuity theorem that applies to more general classes of non necessarily symmetric transition potentials and log-concave type target marginals, including strongly convex at infinity potentials.
The article [eckstein2025hilbert] is also based on a novel version of Hilbert’s projective metric for spaces of integrable functions of bounded growth. Extending results in [chen, deligiannidis] to unbounded cost functions, the recent article  [eckstein2025hilbert] presents exponential decays of these new classes of Hilbert’s projective metrics with respect to some cone of functions in settings where the marginal distributions have sufficiently light tails compared to the growth of the cost function. These results also imply the exponential convergence of the total variation distance. On normed finite dimensional spaces, the main condition in  [eckstein2025hilbert] is that $W$ grows faster that a polynomial of order $p$ and the tails of both marginals decay faster that $e^{-r^{p+q}}$ for some $q>0$. This condition doesn’t cover Gaussian models.
@@
As underlined in [adm-24], Sinkhorn semigroups belong to the class of time varying Markov processes sharing a common invariant target measure at each individual step. These classes of time varying Markov semigroups arise in a variety of areas in applied mathematics including on nonlinear filtering, physics and molecular chemistry, see for instance [dpa, dm-04, dg-01, horton-dp, saloff-zuniga, saloff-zuniga-2, saloff-zuniga-3] and references therein.
The stability analysis of time inhomogenous Markov processes is a broad and notoriously difficult subject, much more difficult to handle than their time homogeneous counterparts; this is because the operators may drastically change during the semigroup evolution.
@@
The articles [douc, saloff-zuniga, saloff-zuniga-2] provide several successful probabilistic and analytical tools, including coupling methods, spectral analysis as well as functional inequalities such as Nash or log-Sobolev inequalities. Sinkhorn semigroups being defined sequentially by nonlinear conjugate transformations, the application of these refined analytical techniques in this context is rather challenging. There is a need for new tools in order to handle these models. In this paper, we introduce a novel approach to Sinkhorn stability analysis that relies on Lyapunov techniques and contraction coefficients on weighted Banach spaces.
1.3 Contributions
@@
We investigate Lyapunov approaches and related weighted-norm contraction techniques to analyze the convergence of Sinkhorn semigroups. These semigroup techniques are basic quantitative methods for studying the stability of nonlinear and time-varying stochastic models, see for instance [horton-dp] as well as Section 2 in [dpa]. In the context of time-homogeneous models, we also refer to the pioneering articles by Hasminskii [hasminskii] and Meyn and Tweedie [meyn-1], see also the book [meyn-2] and references therein.
These powerful off-the-shelf semigroup techniques are mainly based on two conditions: 1) the existence of a Lyapunov function satisfying a drift-type condition; 2) a local minorization condition on the sub-level sets of the Lyapunov function.
@@
Such drift-minorization approaches are well-established techniques to localize the minorization conditions on the compact sub-level sets of the Lyapunov function. In this context, Markov operators act on weighted Banach spaces and operator norms coincide with the contraction coefficients of the Markov integral operator. For a more detailed discussion on this contraction coefficients we refer to [dpa, horton-dp, penev].
This Lyapunov approach and the operator-theoretic framework presented in this article strongly differ from the approaches discussed earlier and taken in the series of recent articles [chiarini, durmus, del2025stability, eckstein2025hilbert].
Our main contributions can be summarized in the following way:
  • In the context of bounded transition potentials, we provide novel strict contraction estimates with respect to general $\Phi$-divergence criteria, including $\mathbb{L}_{p}$-norms, Havrda-Charvat entropies, $\alpha$-divergences, Jensen-Shannon divergence, squared Hellinger integrals and Kakutani-Hellinger, as well as Jeffreys and Rényi divergences [csiszar, dm-03, majernik, nielsen] (see, for instance, Theorem 2.2 and Section 4.1).
  • In the context of locally bounded cost functions and marginal potentials with compact sub-level sets, we develop new weighted total variation norm estimates, including Kantorovich and Wasserstein-type distances (see Section 4.4 and Section 4.5) as well as relative entropy bridge estimates (in Section 4.6).
    We also emphasize that the choice of the Wasserstein semi-metric and the Lyapunov functions depends on the marginal target measures (see Theorem 2.3 and Section 4.5).
  • We develop general Lyapunov contractions principles under practical and minimal regularity conditions that allow to provide quantitative exponential stability estimates for a large class of models.
    We apply this novel framework in a variety of situations, ranging from polynomial growth potentials and heavy tailed marginals on general normed spaces to more sophisticated boundary state space models, including semi-circle transitions, Beta, Weibull, exponential marginals as well as semi-compact models (e.g., semi-discrete marginal models). Last but not least, our approach also allows to consider statistical finite mixture of the above models including kernel-type density estimators of complex data distributions arising in generative models.
To the best of our knowledge, the stability results discussed above are the first results of this type in the literature on entropic optimal transport and Sinkhorn semigroups.
1.4 Outline of the article
@@
After presenting some basic notation in Section 1.5, we begin with a brief review on Schrödinger bridge problems and Sinkhorn semigroups in Section 2 (and we refer to [adm-24] for a more detailed discussion on these models within our framework).
Section 2.3, presents a brief summary of the main stability theorems of the article. All the methodologies and regularity conditions discussed in the article are illustrated in a variety of examples in Section 5. To motivate our results, however, Section 2.4 already includes the discussion of some application models. We also give some detailed comments on our regularity conditions and propose an avenue of open research projects.
Some background on Lyapunov approaches and weighted-norm contraction properties of Markov integral operators is provided in Section 3, with additional details in Appendix C. Section 3.1 recalls some basics facts on the Hilbert projective metric, operator norms, entropy and related $\phi$-divergences. The operator contraction coefficients associated with these criteria are introduced in Section 3.2.
@@
With these notations at hand, Section 4 shows how to adapt this operator-theoretic framework to the convergence analysis of Sinkhorn semigroups.
@@
Section 4.2 presents some general drift-minorization conditions that enable us to estimate the contraction coefficients of Sinkhorn semigroups. The main contraction theorems introduced in this article are presented in Section 4.3. Section 4.4 presents a simple and new regularity condition that allows to provide quantitative exponential stability estimates for a large class of models. Section 4.5 presents quantitative exponential decays w.r.t. Wasserstein semi-metrics and Section 4.6 is dedicated to relative entropy decays and the convergence of Sinkhorn bridges.
Section 5 is dedicated to a variety of illustrations. In particular, polynomial growth models are discussed in Section 5.1, Section 5.2 is dedicated to heavy tailed models, and in Section 5.3 we study bounded Lipschitz domains. Beta and Weibull distributions are discussed, respectively, in Section 5.4 and Section 5.5. An explicit description of the key regularity conditions and results in the context of multivariate linear Gaussian distributions is provided in Section 5.8. Finally, Section 5.9 is dedicated to statistical finite (non necessarily Gaussian) mixture models.
Detailed proofs and some extended analyses can be found in the appendices to this manuscript. Appendix A is devoted to the Gaussian Sinkhorn algorithm, Appendices B and C contain material related to Hilbert metric and weighted norm contractions, respectively. Appendix D is devoted to Schrödinger potential functions. Finally, some additional technical proofs provided for completeness in Appendix E.
@@ 1.5 Some basic notation
Let ${\cal M}(\mathbb{X})$ be the set of locally bounded signed measures on $\mathbb{X}$, and let ${\cal M}_{b}(\mathbb{X})\subset{\cal M}(\mathbb{X})$ the subset of bounded measures. We denote the convex subset of probability measures on $\mathbb{X}$ as ${\cal M}_{1}(\mathbb{X})\subset{\cal M}_{b}(\mathbb{X})$.
The Lebesgue integral of an integrable function $f\in{\cal B}(\mathbb{X})$ w.r.t. $\nu\in{\cal M}(\mathbb{X})$ is denoted by
$\nu(f):=\int_{\mathbb{X}}f(x)~\nu(dx).$
For indicator functions $f=1_{A}$ of a measurable subset $A\subset\mathbb{X}$ sometimes we write $\nu(A)$ instead of $\nu(1_{A})$.
@@
$(\mu{\cal K})(dy):=\int~\mu(dx){\cal K}(x,dy)\quad\mbox{\rm and}\quad{\cal K}(f)(x):=\int{\cal K}(x,dy)f(y).$ (6)
@@
Given a Markov transition ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and a measure $\mu\in{\cal M}(\mathbb{X})$ we set
$(\mu\times{\cal K})(d(x,y)):=\mu(dx)\,{\cal K}(x,dy)\quad\mbox{\rm and}\quad(\mu\times{\cal K})^{\flat}(d(y,x)):=\mu(dx)\,{\cal K}(x,dy).$
@@
The positive and negative part of $a\in\mathbb{R}$ are defined respectively by $a_{+}=\max(a,0)$ and $a_{-}=\max(-a,0)$. Given $a,b\in\mathbb{R}$ we also set $a\vee b=\max(a,b)$ and $a\wedge b=\min(a,b)$.
Kantorovich and Wasserstein criteria
We consider the weighted total variation norm associated with some positive function $\varphi\in{\cal B}(\mathbb{X})$ defined as
[…]
where $\Pi(\mu_{1},\mu_{2})$ stands for the set of coupling probability measures $\cal P$} on the product space $(\mathbb{X}\times\mathbb{X})$, with prescribed first and second coordinate marginals $(\mu_{1},\mu_{2})$ such that $\mu_{i}(\varphi)<\infty$. For a given semi-metric $\varrho(x_{1},x_{2})$ on $\mathbb{X}$ and $p\geq 1$ we also consider the Wasserstein semi-metric $\mathbb{W}_{p,\varrho}$ defined by
$\mathbb{W}_{p,\varrho}(\mu_{1},\mu_{2}):=\inf_{{\cal P}\in\Pi(\mu_{1},\mu_{2})}\left(\int~\varrho(x_{1},x_{2})^{p}~{\cal P}(d(x_{1},x_{2}))\right)^{1/p}.$ (8)
[…]
When $\varrho(x_{1},x_{2})$ is a metric on $\mathbb{X}$ (possibly different from the metric of the Polish space) we recall (see for instance [edwards2011kantorovich, fernique-81, kellerer1982duality, Szulga-83, villani2008optimal]) the Kantorovich-Rubinstein duality
$\mathbb{W}_{\varrho}(\mu_{1},\mu_{2})=\sup{\left\{(\mu_{1}-\mu_{2})(f)~;~f~:~\mathbb{X}\mapsto\mathbb{R}\quad s.t.\quad|f(x_{1})-f(x_{2})|\leq\varrho(x_{1},x_{2})\right\}}.$
Note that
[…]
More generally, for any given semi-metric $\varrho(x_{1},x_{2})$ on $\mathbb{X}$ such that
$\varphi(x_{1})+\varphi(x_{2})\geq\varrho(x_{1},x_{2})\quad\mbox{\rm and}\quad\varphi(x)\geq c>0$ (9)
we have the estimates
[…]
For instance on a given normed vector space $(\mathbb{X},\|\mbox{\LARGE.}\|)$, for any $b,c>0$ and $p,q\geq 1$, if
$\varphi(x):=c~\exp{\left(b\|x\|^{p}\right)}$
then
$\varphi(x_{1})+\varphi(x_{2})\geq\varrho_{\tiny exp}(x_{1},x_{2}):=2c~\left(\exp{\left(\frac{b}{2^{1+(p-1)_{+}}}~\|x_{1}-x_{2}\|^{p}\right)}-1\right).$
We check this claim using the fact that $u+v\geq 2\sqrt{u}\sqrt{v}$ and $(u^{p}+v^{p})\geq 2^{-(p-1)_{+}}(u+v)^{p}$ for any $p>0$ and $u,v\geq 0$. In this context, we have
[…]
The choice of the semi-metric $\varrho$ in (10) is not unique. Indeed, developing the exponential function for any $q\geq 1$ we also have the estimate
$\varphi(x_{1})+\varphi(x_{2})\geq\varrho_{p,q}(x_{1},x_{2}):=c_{p,q}~\|x_{1}-x_{2}\|^{pq}\quad\mbox{\rm with}\quad c_{p,q}:=2c~\frac{1}{q!}\left(\frac{b}{2^{1+(p-1)_{+}}}\right)^{q},$ (11)
and
[…]
Entropy and divergences
Consider a convex function
$\begin{array}[]{cccl}\Phi:&\mathbb{R}_{+}^{2}&\mapsto&\mathbb{R}\cup\{+\infty\}\\ &(u,v)&\leadsto&\Phi(u,v)\quad\mbox{\rm s.t. $\forall a\in\mathbb{R}_{+}$~~~ $\Phi(au,av)=a~\Phi(u,v)$ and $\Phi(1,1)=0$.}\end{array}$ (12)
$D_{\Phi}(\nu_{1},\nu_{2}):=\int~\Phi\left(\frac{d\nu_{1}}{d\gamma}(x),\frac{d\nu_{2}}{d\gamma}(x)\right)~\gamma(dx).$ (13)
By homogeneity arguments, the definition (13) does not depend on the choice of the dominating measure $\gamma$. Note that
Whenever $\Phi(1,0)=\infty$, the $\Phi$-entropy coincides with the $\phi$-divergence in the sense of Csiszár [csiszar], i.e., we also have
$\Phi(u,v)=\Phi_{1}(u,v):=u\log{(u/v)}\quad\mbox{\rm and}\quad\phi(u)=u\log{u}\$
we recover the classical formulae
$\phi_{1}(u)=u\log{u}-(u-1).$
2 Statement of main results
2.1 Sinkhorn bridges
The Schrödinger bridge distribution $\mathbb{P}$ defined in (4) can rarely be computed analytically. However, solutions can be approximated efficiently using the Sinkhorn algorithm, also referred to as the iterative proportional fitting procedure [cuturi, sinkhorn-2, sinkhorn-3]. This procedure is defined sequentially for any $n\geq 0$ by a collection of probability distributions (${\cal P}_{n}$) and transition operators (${\cal S}_{n}$) given by